OpenStudy (anonymous):

dv=e^6x dx v=ind_int e^6x dx=1/6 e^6x Can someone kindly explain this?

OpenStudy (anonymous):

Is that \[e^6x\] or\[e^{6x}\]

OpenStudy (anonymous):

could someone help me with my problem...im about to bump it in a minute :)

OpenStudy (anonymous):

e^6x

OpenStudy (anonymous):

I know that ind_int e^x= e^x...it's the 1/6 I'm confused about!

OpenStudy (anonymous):

My browser crashed, have to type again.. sigh..

OpenStudy (anonymous):

Thanks :-)

OpenStudy (anonymous):

\[\int\limits{}e^{6x}dx\] Use u-substitution \[u=e^{6x}\]\[du=6e^{6x}\] \[\frac{1}{6}\int\limits{6e^{6x}}dx=\frac{1}{6}\int\limits{}udu\] Can you solve from here?

OpenStudy (anonymous):

I'll give it a try, thanks

OpenStudy (anonymous):

Hold on typo Use\[u=6x\]\[du=6\] \[\frac{1}{6}\int\limits{}6e^{6x}=\frac{1}{6}\int\limits{}e^udu\]

OpenStudy (anonymous):

i saw that :-)...I was forgetting to use u-sub within the integration by parts.

OpenStudy (anonymous):

Yea brain farts... Where did you use integration by parts on this probelm?

OpenStudy (anonymous):

\[\int\limits_{}^{}xe ^{6x}dx\]

OpenStudy (anonymous):

Ah I see, the original question was just part of working your previous comment.